I need some help in solving this problem:

2x + 4 = -8
Thanks, Lisa

Answer by funmath(2933)

Hi Lisa,
The goal is to get x all by itself on one side of the equal sign within the rules of algebra. What ever you do to one side of the = siqn you must also do to the other.
2x + 4 = -8 The 4 doesn't have an x so we need to substract it from both sides.
2x+4-4=-8-4
2x+0=-12
2x=-12 The x is being multiplied by 2, so divide both sides by 2.
2x/2=-12/2
1x=-6
x=-6
:
Check by substituting x=-6 into the original equation and see if both sides are equal after we simplify:
2(-6)+4=-8
-12+4=-8
-8=-8
We're right!!!
Happy Calculating, Lisa!!!